A compound pendulum of mass 0.8 kg and length 1.2 m swings about a pivot point located at one end of the pendulum.


(a) Find its frequency of oscillation.


(b) Find the center of oscillation of the pendulum.

1
Expert's answer
2022-01-04T10:22:48-0500

(a) Frequency:


"f=\dfrac 1T=\dfrac1{2\pi}\sqrt{\dfrac{mgd}{I}}=\dfrac1{2\pi}\sqrt{\dfrac{mgd}{md^2}}=0.45\text{ Hz}."


(b) The center of oscillation of the pendulum 1.2 m above the mass.